OpenStudy (anonymous):

Find the equation in standard form which passes through (–2, –7) and (1, –1) having a brainfart, need a little help. /:

OpenStudy (anonymous):

(-7+1)/(-2-1)=-6/-3=2=m y=mx+b -1=2*1+b -1=2+b -3=b y=2x-3

OpenStudy (anonymous):

test with other points to make sure, -7=-2*2-3 -7=-4-3 -7=-7 yup ^^